OpenStudy (anonymous):

giving medal Write a rule for the linear function whose graph has slope m and contains the given point. m =-1/4; (4, 0) a. y = -1/4x + 1 c. y =1/4x - 4 b. y = -1/4x+ 4 d. y = -1/4x- 1

OpenStudy (souvik):

\[y-y _{1}=m(x-x _{1})\] here m=slope=-1/4

OpenStudy (anonymous):

ya y=-1/4 (x-4)